# On the Robustness of Kernel Goodness-of-Fit Tests

### v4

@@ A.6.1 Statement and Proof of Lemma˜21
Proof of Lemma˜21  The assumption $\mathbb{E}_{\mathbf{Y}\sim R}[u_{p}(\mathbf{Y},\mathbf{Y})^{1/2}]<\infty$ implies that $D(R,P)$ is well-defined by Gorham and Mackey, (2017, Proposition 2). Moreover, when $k\in\mathcal{C}^{(1,1)}_{b}$ and $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$, Gorham and Mackey, (2015, Proposition 1) shows that $\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\cdot)]=0$. Using this and the linearity of expectation,
@@ A.6.2 Proof of Proposition˜5
The proof follows a similar approach as Chérief-Abdellatif and Alquier, (2022, Lemma 3.3). Under the assumed kernel conditions, Lemma˜2 shows that $\sup_{\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x}^{\prime})\leq\tau_{\infty}<\infty$, which implies $D^{2}(Q,P)=\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim Q}[u_{p}(\mathbf{X},\mathbf{X}^{\prime})]\leq\tau_{\infty}<\infty$ for any probability measure $Q\in\mathcal{P}(\mathbb{R}^{d})$. In particular, $D(Q,P)$ is well-defined for any $Q\in\mathcal{P}(\mathbb{R}^{d})$ and the conditions in Lemma˜21 are met. For any $Q=(1-\epsilon)P+\epsilon R$ with $Q\in\mathcal{P}(P;\epsilon_{0})$, applying Lemma˜21 shows that
@@ A.6.4 KSD Balls and Fat Tails
Proof of Proposition˜22  For $\nu>2$, the change-of-variable formula shows that the scaled t-distribution $Q_{\nu}$ has density function $q(x)=Z_{\nu}q_{\nu}^{\ast}(x)$, where $Z_{\nu}\coloneqq\Gamma(\frac{\nu+1}{2})/\big{(}\sqrt{\pi(\nu-2)}\Gamma(\frac{\nu}{2})\big{)}$, $q_{\nu}^{\ast}(x)\coloneqq(1+\frac{x^{2}}{\nu-2})^{-(\nu+1)/2}$, and $\Gamma(\cdot)$ is the Gamma function (see, e.g., Forbes et al.,, 2011, Chapter 8.1). To show $q_{\nu}(x)$ intersects with $p(x)$ at exactly two points on $(0,\infty)$, we first show that $q_{\nu}(x)$ is strong super-Gaussian (Palmer et al.,, 2010), i.e., $x\mapsto\log q_{\nu}(\sqrt{x})$ is convex on $0,\infty)$. This is immediate by writing
@@
and noting that $u\mapsto\log(1+u)$ is concave on $0,\infty)$. Palmer et al., (2010, Theorem 2) shows that any symmetric and strongly super-Gaussian density belongs to the class of density cross inequalities ($DC_{+}$), which are symmetric densities that cross a Gaussian density of equal variance exactly four times on $\mathbb{R}$, and take higher values than that normal density at $x=0$ and as $x\to\infty$. Since in this case both $Q_{\nu}$ and $P$ have unit variance, we conclude that $q_{\nu}$ and $p$ intersect at exactly four points on $\mathbb{R}$. By symmetry of these densities about $x=0$, this shows that $q_{\nu}$ and $p$ must intersect at exactly two points on $(0,\infty)$.
@@ B.1 Decay Rate of Contamination Ratio
Figure 9: Rejection probability for an IMQ kernel with bandwidth $\lambda$. “med” is the median heuristic. “KSDAgg” is the test of Schrab et al., (2022). The dashed line shows $\alpha=0.05$. All tests all reject $H_{0}$ for large $z$.
@@ B.2 Implementation Details for the Robust MMD Tests
This section provides implementation details for the two robust MMD-based tests included in Section˜5.1. Both tests use the Maximum Mean Discrepancy (MMD; Gretton et al.,, 2012, Definition 2) between the data-generating distribution $Q$ and the model $P$ as a measure of their disparity. Given a reproducing kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$, the squared MMD between $Q$ and $P$ are defined as
@@
Given independent samples $\mathbb{X}_{n}=\{\mathbf{X}_{i}\}_{i=1}^{n}$ from $Q$ and $\mathbb{Y}_{m}=\{\mathbf{Y}_{j}\}_{j=1}^{m}$ from $P$, the MMD can be estimated directly by $D_{\mathrm{MMD}}^{2}(Q_{n},P_{m})$, where $Q_{n}$ and $P_{m}$ are the empirical measures formed by $\mathbb{X}_{n}$ and $\mathbb{Y}_{m}$, respectively (Gretton et al.,, 2012). The computational cost of such estimate scales with sample size $n$ as $\mathcal{O}((n+m)^{2}+mC_{\text{sim}})$, where $C_{\text{sim}}$ is the cost of simulating one datum from $P$, compared with $\mathcal{O}(n^{2}+nC_{\text{score}})$ for KSD, where $C_{\text{score}}$ denotes the cost of a single score evaluation. In our experiment in Section˜5.1, $P$ is a Gaussian model, so both simulation and score evaluation are trivial and $C_{\text{sim}},C_{\text{score}}$ are both small. Hence, we set $m=n$ so that the costs of MMD and of KSD are comparable. Moreover, we will use the squared-exponential kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\exp(-\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}/(2\gamma^{2}))$ and set $\gamma^{2}$ using the median heuristic. Squared-exponential kernels are popular for MMDs and are also used in the original papers that proposed the two robust MMD tests (Sun and Zou,, 2023; Schrab and Kim,, 2024).
@@ B.2.1 The MMD-Dev Test
The robust MMD test of Sun and Zou, (2023, Eq. (38)), which we refer to as MMD-Dev, targets the hypotheses
@@
where $\theta_{\mathrm{MMD}}\geq 0$ and $\mathcal{B}^{\mathrm{MMD}}(P;\theta_{\mathrm{MMD}})=\{Q\in\mathcal{P}:D_{\mathrm{MMD}}(Q,P)\leq\theta_{\mathrm{MMD}}\}$ is the MMD ball centered at $P$ and with radius $\theta_{\mathrm{MMD}}$. Given a test level $\alpha\in(0,1)$, MMD-Dev rejects $H_{0}^{\mathrm{MMD}}$ if $D_{\mathrm{MMD}}(Q_{n},P_{n})>\theta_{\mathrm{MMD}}+\gamma_{n}$, where $\gamma_{n}=\sqrt{2K/n}(1+\sqrt{-\log\alpha})$. Sun and Zou, (2023, Theorem 4) shows that this test controls the Type-I error for any finite $n$, and is asymptotically optimal against certain alternatives. The decision threshold $\gamma_{n}$ is derived using a McDiarmid-type deviation inequality for MMDs due to Gretton et al., (2012, Theorem 8).
@@
Proof of Lemma˜23  For any $Q\in\mathcal{P}(\mathbb{R}^{d})$, we define its kernel mean embedding (Berlinet and Thomas-Agnan,, 2004, Chapter 4) as $\mu_{Q}(\cdot)\coloneqq\mathbb{E}_{\mathbf{X}\sim Q}[k(\mathbf{X},\cdot)]$, which is well-defined since $k$ is bounded. By Sriperumbudur et al., (2010, Theorem 1), the MMD between any $Q,P\in\mathcal{P}(\mathbb{R}^{d})$ can be equivalently written as $D_{\mathrm{MMD}}(Q,P)=\|\mu_{Q}-\mu_{P}\|_{\mathcal{H}_{k}}$, where $\mathcal{H}_{k}$ is the RKHS associated with $k$. Pick $\epsilon_{0}\in[0,1]$ and $R\in\mathcal{P}$. For any $\epsilon\in[0,\epsilon_{0}]$,
@@ B.2.2 The dcMMD Test
The dcMMD test (Schrab and Kim,, 2024) targets different hypotheses. Given $\epsilon_{0}\in[0,1]$, the null assumes that the observed sample $\mathbb{X}_{n}$ is generated by firstly drawing i.i.d. random variables from some probability measure, then corrupting them by replacing at most a proportion of $\epsilon_{0}$ of the sample by arbitrary values. The hypotheses can be formalized as follows:
@@
The dcMMD test rejects $H_{0}^{\mathrm{dcMMD}}$ if $D_{\mathrm{MMD}}(Q_{n},P_{n})>q_{\alpha}^{\mathrm{MMD}}+2\epsilon_{0}\sqrt{2K}$, where $q_{\alpha}^{\mathrm{MMD}}$ is the empirical quantile of $B$ permutation samples $\{T_{\mathrm{MMD}}^{b}\}_{b=1}^{B}$, and each $T_{\mathrm{MMD}}^{b}$ is computed by (i) randomly permuting $\mathbb{X}_{n}\cup\mathbb{Y}_{n}$, (ii) partitioning the permuted set into two subsets $\mathbb{X}_{n}^{b}$ and $\mathbb{Y}_{n}^{b}$ of size $n$, and (iii) computing $T_{\mathrm{MMD}}^{b}\coloneqq D_{\mathrm{MMD}}(Q_{n}^{b},P_{n}^{b})$, where $Q_{n}^{b}$ and $P_{n}^{b}$ are the empirical measures based on $\mathbb{X}_{n}^{b}$ and $\mathbb{Y}_{n}^{b}$, respectively. We use $B=500$, which is the default setup in Schrab and Kim, (2024).
@@ B.6 Scalability with Dimension
We run the Gaussian mean-shift experiment in Section˜B.5 in different dimensions; other experimental setups remain unchanged. The uncertainty radius for R-KSD is chosen to be the (non-squared) KSD value corresponding to mean-shift $\mu=0.3$. Results are reported in Figure˜12. As dimension increases, both the two standard tests and R-KSD have declining power in correctly rejecting for large values of $\mu$. This shows that both the standard and the robust KSD tests suffer from the curse-of-dimensionality, a known issue for kernel-based tests (Huang et al.,, 2023; Reddi et al.,, 2015; Ramdas et al.,, 2015). Unsurprisingly, this issue is more prominent for the R-KSD test. This is because the KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ could potentially include more distributions in higher dimensions.
@@ B.8 Weighted Bootstrap and Wild Bootstrap
We compare the robust-KSD test using the weighted bootstrap described in Section˜2.2 (equivalent to the Efron’s bootstrap) against the wild bootstrap due to Leucht and Neumann, (2013). For the standard KSD test, weighted bootstrap was used in Liu et al., (2016), while the wild bootstrap is more popular in the literature (Chwialkowski et al.,, 2016; Schrab et al.,, 2022; Liu et al.,, 2023). Compared with weighted bootstrap, the wild approach is more flexible as it can work for dependent samples (Chwialkowski et al.,, 2014), and theoretical guarantees on the power of the standard test were only proved with the wild bootstrap (Schrab et al.,, 2022). In this work, we have used the weighted approach since we can then leverage existing theoretical results to show the validity of the test; see Section˜A.5.1. This is for simplicity rather than necessity.
@@ Appendix C Extension to U-statistics
In this work, we have focused primarily on using the V-statistic (2) for estimating the KSD. An alternative estimator commonly used in the literature (Liu et al.,, 2016; Jitkrittum et al.,, 2017; Schrab et al.,, 2022; Liu et al.,, 2023) is the following U-statistic $\frac{1}{n(n-1)}\sum_{1\leq i\neq j\leq n}u_{p}(\mathbf{X}_{i},\mathbf{X}_{j})$. In this section, we discuss briefly how to extend our key results to this U-statistic estimate.
  • Extending the qualitative (non-)robustness results. The main ingredients for proving Theorem˜1 are (i) a deviation bound for the V-statistic $D^{2}(\mathbb{X}_{n})$ as shown in Lemma˜7, and (ii) a concentration bound for the bootstrapped quantile $q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n})$ as shown in Lemma˜8. A U-statistic counterpart of both results can be shown by following the same proof technique and replacing $D^{2}(\mathbb{X}_{n})$ with the U-statistic. For example, with the U-statistic, a decomposition similar to (10) can be derived by summing over only the non-diagonal terms $i\neq j$ and replacing normalizing factors of the form $1/n^{2}$ to 1/(n(n1))1/(n(n-1)). Similarly, the proof of Theorem˜3 can be extended to U-statistics by following the same steps in Section˜A.3.
  • Extending the robust-KSD test. To adapt the robust-KSD test to the U-statistic, we first note that rejecting the null if $\Delta_{\theta}(\mathbb{X}_{n})>q_{\infty,1-\alpha}$ is equivalent to rejecting the null if $D^{2}(\mathbb{X}_{n})>(q_{\infty,1-\alpha}+\theta)^{2}$, as argued in (53) and the paragraph thereafter. A natural approach is therefore to replace the V-statistic $D^{2}(\mathbb{X}_{n})$ by the U-statistic, and the bootstrap quantile $q_{\infty,1-\alpha}$ by the square-root of the bootstrap quantile formed with the U-statistic. However, since the bootstrap samples (3) based on U-statistics can take negative values, the bootstrap quantile can also take negative values, rendering its square-root $q_{\infty,1-\alpha}$ undefined. One solution is to never reject the null when this happens. This will make little impact when $D^{2}(Q,P)$ is large so that the bootstrap samples very likely take positive values, but will make the test conservative when $D^{2}(Q,P)\approx 0$ so that the bootstrap samples can take negative values with high probability.

### v5

@@ A.6.1 Statement and Proof of Lemma˜21
Proof of Lemma˜21 The assumption $\mathbb{E}_{\mathbf{Y}\sim R}[u_{p}(\mathbf{Y},\mathbf{Y})^{1/2}]<\infty$ implies that $D(R,P)$ is well-defined by Gorham and Mackey (2017, Proposition 2). Moreover, when $k\in\mathcal{C}^{(1,1)}_{b}$ and $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$, Gorham and Mackey (2015, Proposition 1) shows that $\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\cdot)]=0$. Using this and the linearity of expectation,
@@ A.6.2 Proof of Proposition˜5
Proof The proof follows a similar approach as Chérief-Abdellatif and Alquier (2022, Lemma 3.3). Under the assumed kernel conditions, Lemma˜2 shows that $\sup_{\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x}^{\prime})\leq\tau_{\infty}<\infty$, which implies $D^{2}(Q,P)=\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim Q}[u_{p}(\mathbf{X},\mathbf{X}^{\prime})]\leq\tau_{\infty}<\infty$ for any probability measure $Q\in\mathcal{P}(\mathbb{R}^{d})$. In particular, $D(Q,P)$ is well-defined for any $Q\in\mathcal{P}(\mathbb{R}^{d})$ and the conditions in Lemma˜21 are met. For any $Q=(1-\epsilon)P+\epsilon R$ with $Q\in\mathcal{P}(P;\epsilon_{0})$, applying Lemma˜21 shows that
@@ A.6.4 KSD Balls and Fat Tails
Proof of Proposition˜22 For $\nu>2$, the change-of-variable formula shows that the scaled t-distribution $Q_{\nu}$ has density function $q(x)=Z_{\nu}q_{\nu}^{\ast}(x)$, where $Z_{\nu}\coloneqq\Gamma(\frac{\nu+1}{2})/\big(\sqrt{\pi(\nu-2)}\Gamma(\frac{\nu}{2})\big)$, $q_{\nu}^{\ast}(x)\coloneqq(1+\frac{x^{2}}{\nu-2})^{-(\nu+1)/2}$, and $\Gamma(\cdot)$ is the Gamma function (see, e.g., Forbes et al., 2011, Chapter 8.1). To show $q_{\nu}(x)$ intersects with $p(x)$ at exactly two points on $(0,\infty)$, we first show that $q_{\nu}(x)$ is strong super-Gaussian (Palmer et al., 2010), i.e., $x\mapsto\log q_{\nu}(\sqrt{x})$ is convex on $0,\infty)$. This is immediate by writing
@@
and noting that $u\mapsto\log(1+u)$ is concave on $0,\infty)$. Palmer et al. (2010, Theorem 2) shows that any symmetric and strongly super-Gaussian density belongs to the class of density cross inequalities ($DC_{+}$), which are symmetric densities that cross a Gaussian density of equal variance exactly four times on $\mathbb{R}$, and take higher values than that normal density at $x=0$ and as $x\to\infty$. Since in this case both $Q_{\nu}$ and $P$ have unit variance, we conclude that $q_{\nu}$ and $p$ intersect at exactly four points on $\mathbb{R}$. By symmetry of these densities about $x=0$, this shows that $q_{\nu}$ and $p$ must intersect at exactly two points on $(0,\infty)$.
@@ B.1 Decay Rate of Contamination Ratio
Figure 9: Rejection probability for an IMQ kernel with bandwidth $\lambda$. “med” is the median heuristic. “KSDAgg” is the test of Schrab et al. (2022). The dashed line shows $\alpha=0.05$. All tests all reject $H_{0}$ for large $z$.
@@ B.2 Implementation Details for the Robust MMD Tests
This section provides implementation details for the two robust MMD-based tests included in Section˜5.1. Both tests use the Maximum Mean Discrepancy (MMD; Gretton et al., 2012, Definition 2) between the data-generating distribution $Q$ and the model $P$ as a measure of their disparity. Given a reproducing kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$, the squared MMD between $Q$ and $P$ are defined as
@@
Given independent samples $\mathbb{X}_{n}=\{\mathbf{X}_{i}\}_{i=1}^{n}$ from $Q$ and $\mathbb{Y}_{m}=\{\mathbf{Y}_{j}\}_{j=1}^{m}$ from $P$, the MMD can be estimated directly by $D_{\mathrm{MMD}}^{2}(Q_{n},P_{m})$, where $Q_{n}$ and $P_{m}$ are the empirical measures formed by $\mathbb{X}_{n}$ and $\mathbb{Y}_{m}$, respectively (Gretton et al., 2012). The computational cost of such estimate scales with sample size $n$ as $\mathcal{O}((n+m)^{2}+mC_{\text{sim}})$, where $C_{\text{sim}}$ is the cost of simulating one datum from $P$, compared with $\mathcal{O}(n^{2}+nC_{\text{score}})$ for KSD, where $C_{\text{score}}$ denotes the cost of a single score evaluation. In our experiment in Section˜5.1, $P$ is a Gaussian model, so both simulation and score evaluation are trivial and $C_{\text{sim}},C_{\text{score}}$ are both small. Hence, we set $m=n$ so that the costs of MMD and of KSD are comparable. Moreover, we will use the squared-exponential kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\exp(-\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}/(2\gamma^{2}))$ and set $\gamma^{2}$ using the median heuristic. Squared-exponential kernels are popular for MMDs and are also used in the original papers that proposed the two robust MMD tests (Sun and Zou, 2023; Schrab and Kim, 2024).
@@ B.2.1 The MMD-Dev Test
The robust MMD test of Sun and Zou (2023, Eq. 38), which we refer to as MMD-Dev, targets the hypotheses
@@
where $\theta_{\mathrm{MMD}}\geq 0$ and $\mathcal{B}^{\mathrm{MMD}}(P;\theta_{\mathrm{MMD}})=\{Q\in\mathcal{P}:D_{\mathrm{MMD}}(Q,P)\leq\theta_{\mathrm{MMD}}\}$ is the MMD ball centered at $P$ and with radius $\theta_{\mathrm{MMD}}$. Given a test level $\alpha\in(0,1)$, MMD-Dev rejects $H_{0}^{\mathrm{MMD}}$ if $D_{\mathrm{MMD}}(Q_{n},P_{n})>\theta_{\mathrm{MMD}}+\gamma_{n}$, where $\gamma_{n}=\sqrt{2K/n}(1+\sqrt{-\log\alpha})$. Sun and Zou (2023, Theorem 4) shows that this test controls the Type-I error for any finite $n$, and is asymptotically optimal against certain alternatives. The decision threshold $\gamma_{n}$ is derived using a McDiarmid-type deviation inequality for MMDs due to Gretton et al. (2012, Theorem 8).
@@
Proof of Lemma˜23 For any $Q\in\mathcal{P}(\mathbb{R}^{d})$, we define its kernel mean embedding (Berlinet and Thomas-Agnan, 2004, Chapter 4) as $\mu_{Q}(\cdot)\coloneqq\mathbb{E}_{\mathbf{X}\sim Q}[k(\mathbf{X},\cdot)]$, which is well-defined since $k$ is bounded. By Sriperumbudur et al. (2010, Theorem 1), the MMD between any $Q,P\in\mathcal{P}(\mathbb{R}^{d})$ can be equivalently written as $D_{\mathrm{MMD}}(Q,P)=\|\mu_{Q}-\mu_{P}\|_{\mathcal{H}_{k}}$, where $\mathcal{H}_{k}$ is the RKHS associated with $k$. Pick $\epsilon_{0}\in[0,1]$ and $R\in\mathcal{P}$. For any $\epsilon\in[0,\epsilon_{0}]$,
@@ B.2.2 The dcMMD Test
The dcMMD test (Schrab and Kim, 2024) targets different hypotheses. Given $\epsilon_{0}\in[0,1]$, the null assumes that the observed sample $\mathbb{X}_{n}$ is generated by firstly drawing i.i.d. random variables from some probability measure, then corrupting them by replacing at most a proportion of $\epsilon_{0}$ of the sample by arbitrary values. The hypotheses can be formalized as follows:
@@
The dcMMD test rejects $H_{0}^{\mathrm{dcMMD}}$ if $D_{\mathrm{MMD}}(Q_{n},P_{n})>q_{\alpha}^{\mathrm{MMD}}+2\epsilon_{0}\sqrt{2K}$, where $q_{\alpha}^{\mathrm{MMD}}$ is the empirical quantile of $B$ permutation samples $\{T_{\mathrm{MMD}}^{b}\}_{b=1}^{B}$, and each $T_{\mathrm{MMD}}^{b}$ is computed by (i) randomly permuting $\mathbb{X}_{n}\cup\mathbb{Y}_{n}$, (ii) partitioning the permuted set into two subsets $\mathbb{X}_{n}^{b}$ and $\mathbb{Y}_{n}^{b}$ of size $n$, and (iii) computing $T_{\mathrm{MMD}}^{b}\coloneqq D_{\mathrm{MMD}}(Q_{n}^{b},P_{n}^{b})$, where $Q_{n}^{b}$ and $P_{n}^{b}$ are the empirical measures based on $\mathbb{X}_{n}^{b}$ and $\mathbb{Y}_{n}^{b}$, respectively. We use $B=500$, which is the default setup in Schrab and Kim (2024).
@@ B.6 Scalability with Dimension
We run the Gaussian mean-shift experiment in Section˜B.5 in different dimensions; other experimental setups remain unchanged. The uncertainty radius for R-KSD is chosen to be the (non-squared) KSD value corresponding to mean-shift $\mu=0.3$. Results are reported in Figure˜12. As dimension increases, both the two standard tests and R-KSD have declining power in correctly rejecting for large values of $\mu$. This shows that both the standard and the robust KSD tests suffer from the curse-of-dimensionality, a known issue for kernel-based tests (Huang et al., 2023; Reddi et al., 2015; Ramdas et al., 2015). Unsurprisingly, this issue is more prominent for the R-KSD test. This is because the KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ could potentially include more distributions in higher dimensions.
@@ B.8 Weighted Bootstrap and Wild Bootstrap
We compare the robust-KSD test using the weighted bootstrap described in Section˜2.2 (equivalent to the Efron’s bootstrap) against the wild bootstrap due to Leucht and Neumann (2013). For the standard KSD test, weighted bootstrap was used in Liu et al. (2016), while the wild bootstrap is more popular in the literature (Chwialkowski et al., 2016; Schrab et al., 2022; Liu et al., 2023). Compared with weighted bootstrap, the wild approach is more flexible as it can work for dependent samples (Chwialkowski et al., 2014), and theoretical guarantees on the power of the standard test were only proved with the wild bootstrap (Schrab et al., 2022). In this work, we have used the weighted approach since we can then leverage existing theoretical results to show the validity of the test; see Section˜A.5.1. This is for simplicity rather than necessity.
@@ Appendix C Extension to U-statistics
In this work, we have focused primarily on using the V-statistic (2) for estimating the KSD. An alternative estimator commonly used in the literature (Liu et al., 2016; Jitkrittum et al., 2017; Schrab et al., 2022; Liu et al., 2023) is the following U-statistic $\frac{1}{n(n-1)}\sum_{1\leq i\neq j\leq n}u_{p}(\mathbf{X}_{i},\mathbf{X}_{j})$. In this section, we discuss briefly how to extend our key results to this U-statistic estimate.
  • Extending the qualitative (non-)robustness results: The main ingredients for proving Theorem˜1 are (i) a deviation bound for the V-statistic $D^{2}(\mathbb{X}_{n})$ as shown in Lemma˜7, and (ii) a concentration bound for the bootstrapped quantile $q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n})$ as shown in Lemma˜8. A U-statistic counterpart of both results can be shown by following the same proof technique and replacing $D^{2}(\mathbb{X}_{n})$ with the U-statistic. For example, with the U-statistic, a decomposition similar to (10) can be derived by summing over only the non-diagonal terms $i\neq j$ and replacing normalizing factors of the form $1/n^{2}$ to 1/(n(n1))1/(n(n-1)). Similarly, the proof of Theorem˜3 can be extended to U-statistics by following the same steps in Section˜A.3.
  • Extending the robust-KSD test: To adapt the robust-KSD test to the U-statistic, we first note that rejecting the null if $\Delta_{\theta}(\mathbb{X}_{n})>q_{\infty,1-\alpha}$ is equivalent to rejecting the null if $D^{2}(\mathbb{X}_{n})>(q_{\infty,1-\alpha}+\theta)^{2}$, as argued in (53) and the paragraph thereafter. A natural approach is therefore to replace the V-statistic $D^{2}(\mathbb{X}_{n})$ by the U-statistic, and the bootstrap quantile $q_{\infty,1-\alpha}$ by the square-root of the bootstrap quantile formed with the U-statistic. However, since the bootstrap samples (3) based on U-statistics can take negative values, the bootstrap quantile can also take negative values, rendering its square-root $q_{\infty,1-\alpha}$ undefined. One solution is to never reject the null when this happens. This has little impact when $D^{2}(Q,P)$ is large, since the bootstrap samples are then likely to take positive values. However, when $D^{2}(Q,P)\approx 0$, the bootstrap samples are more prone to taking negative values, thus making the test conservative.